OpenStudy (anonymous):

Can someone please help me with this story problem thing............when the justices of the supreme court assemble to go on the bench each day each justice shakes hands with each of the other justices for a total of 36 handshakes. the total number of handshakes "h" possible for "n" people is given by h=n^2-n/2. write and solve an equation to determine the number of justices on the supreme court.

OpenStudy (anonymous):

substitute h = 36.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

36 = n^2-n/2 solve for n

OpenStudy (anonymous):

yeah i get that equation but i dont get how to solve for n

OpenStudy (anonymous):

have you solved quadratic equations before?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

manipulate your equation so that it resembles 0=n^2-n/2-36. then use the quadratic formula\[(-b \pm \sqrt{b^2-4ac}/2a)\] a=coefficient in front of n^2 b=coefficient in front of n and c is the constant. in this case a=1 b=1/2 and c=36

OpenStudy (anonymous):

huh??

OpenStudy (anonymous):

yes, you can use what ridgway said. there are other ways of solving quadratic equations. Just google "solve quadratic equations". there are a number of videos explaining how to solve quadratic equations.

OpenStudy (anonymous):

ive tried that and none of those make sense to me